eatsalad:

Solve the system of equations. x+2y+3z=3 −x+3y+2z=12 x+y+z=−2

eatsalad:

I begin by eliminating x and adding x+2y+3z=3 −x+3y+2z=12


And I get 5y+5z=15

eatsalad:

Im not sure what to do after that

surjithayer:

5y+5z=15 divide by 5 y+z=3 x+y+z=-2 x+3=-2 x=-2-3=-5 x=-5 from first eq. -5+2y+3z==3 2y+3z=-3-5=-8 y+z=3 multiply by 2 2y+2z=6 2y+3z=-8 subtract -z=6+8=14 z=-14 y+z=3 y+(-14)=3 y=3+14=17 x=-5 y=17 z=-14

eatsalad:

i dont understand honestly